# expected information gain SLAM simultaneous localization and mapping

### v1

@@ I Introduction
Robot manipulators in the wild will inevitably come across previously unseen objects and environments. They can benefit from building object models from the limited data available at test time. Such models can be reused across different tasks and systems, and they can make learned manipulation policies more interpretable. In the building of these models, tactile sensing can compliment visual modalities like RGB and depth. Tactile data is robust to darkness, reflections, and occlusions from clutter, the environment, and the object itself. With the advent of inexpensive sensors [1, 2, 3], touch is becoming a more common component of robot manipulation systems [4].
Two challenges inherent to tactile sensing are sparsity and disturbance. By definition, data is only collected about the part of the object in direct physical contact with the sensor, necessitating multiple movements to get a complete picture. Sparsity can be addressed by active exploration, where actions are taken to efficiently maximize some information metric, often captured by maintaining a belief distribution over possible object parameterizations [5, 6, 7, 8, 9, 10]. However, reasoning about information gain is complicated by disturbance: every contact has the potential to move the object unless it is heavy or bolted down. Propagating belief through motion is tricky, informing our first key insight: expected information gain (EIG), by foregoing an explicit belief distribution, is a particularly useful metric for active tactile exploration.
Disturbance affects even the model-building itself, as it necessitates learning the object’s pose trajectory in addition to its geometry. That trajectory needs to adhere to the physical constraints of rigid-body motion and contact, which can lead to difficult loss landscapes with both nearly flat and nearly discontinuous regions. Related work in geometry-learning with a known trajectory [11, 12] has shown how a violation-implicit loss leads to a landscape more amenable to gradient-based optimization. Our second key insight is that this loss works just as well with the trajectory itself as a decision variable.
@@ II-A Model Building and System Identification
Visual modeling of both environment and object geometry is a mature field with the well-studied paradigms of simultaneous localization and mapping (SLAM) and structure from motion (SfM)[13]. Continuous object representations include radiance fields [14] and Gaussian splatting [15]. Recent work learns the geometry [16] and pose [17] of dynamic objects in video, even in contact-rich environments [18]. For assessing the accuracy of our tactile-only system, we compare against ground truth estimates from FoundationPose [17].
Even without active exploration, tactile sensing has been used in recent work to refine models of both the robot and its environment. Non-optical contact sensors can enable more accurate proprioception in the presence of a known environment [19]. Optical tactile sensors create estimates of local surface geometry that can be used to localize the robot with respect to a known object [20] or vice versa [21].
@@ II-B Active Exploration in Robotics
Active exploration has been studied in robotics for decades [22] and involves either the selection of data-gathering actions to maximize information or the curation of data based on information potential. Generally, it includes: (1) some current state of knowledge, (2) an estimate of the information collected by a given action, and (3) an optimization metric quantifying how useful the latter is given the former.
In Bayesian state estimation, the current state can be represented by a belief distribution. Mutual information (MI) [5] as a metric captures the information gained about this belief distribution by observing some other (e.g. measurement likelihood) distribution, but in dynamic environments this can be intractable to compute exactly [6]. Explicit MI calculation can be avoided by using uncertainty instead, e.g., probing areas where the belief distribution has a high variance [7, 8]. One could also use a heuristic to avoid actions similar to those previously taken [10] or a policy trained to take informative actions [9]. For dynamic objects, propagating belief though a dynamics model that is itself uncertain is tricky: the result is inherently multimodal, and the number of required parameters scales exponentially in the number of moments. Multimodal uncertainty can be approximated by simulating an ensemble of possible models [23].
Another approach is to drop the belief distribution and maximize Fisher Information [14], which only requires a point maximum-likelihood estimate (MLE) and is used extensively in optimal experimental design. Recent work uses this approach to refine a simulator from real world data [24] and plan informative trajectories in contact-rich environments [25]. However, maximizing Fisher information alone does not take into account previously observed information over multiple actions. Ergodic exploration [26, 27] can be used to plan a sequence of actions such that, while more time is spent in areas likely to have high information, the entire state space is eventually explored. EIG (see Sec. V), derived from Fisher information, explicitly quantifies how much information an action is likely to obtain relative to previously observed information. It has been used in visual model building for active view selection of static objects [14, 15]. This work takes this procedure and implements it for tactile sensing of dynamic objects.
@@ II-C Tactile Active Exploration
Multiple works have tackled the problem of tactile active exploration. The most common application is the modeling of an unknown surface using poking or sliding contact from an optical sensor [28, 29, 7], as thoroughly collated by a recent survey [30]. It is also common to estimate the geometry of 3d static objects [10, 8]. When learning both pose and geometry, the latter tends to be treated as a classification problem by maintaining a discrete set of possible object geometries [31, 32, 9]. One work simultaneously estimates the pose and geometry of specifically planar objects [33]. To our knowledge, ours is the first work that learns the pose and geometry (from a continuous class of convex shapes) of a 3D rigid body that is free to move during deployment.
@@ IV Learning with a Violation-Implicit Loss
Where $f_{\theta}$ are the discrete-time dynamics as a function of contact impulses $\lambda_{t}$. These can themselves be written as the minimizer of a function $g_{\theta}$ [34] representing other physical constraints (e.g. non-penetration) discussed in more detail below. Under the assumption of Coulomb friction, impulses are constrained to the friction cone $\mathcal{FC}=\{\|\lambda_{t}^{f}\|_{2}\leq\mu\lambda_{t}^{n}\}$, where the total frictional impulse $\|\lambda_{t}^{f}\|_{2}$ is bounded by a constant multiple of the normal impulse $\lambda_{t}^{n}$, sliding in equality and sticking otherwise.
One could optimize this loss function through shooting and differentiable simulation. With parameter $x_{0}$, compute $x_{t}=(f_{\theta})^{t}(x_{0})$ and perform gradient descent through $(f_{\theta})^{t}$ to find $\tilde{x}_{0}$. Then $\tilde{x}_{T}=(f_{\theta})^{T}(\tilde{x}_{0})$. As discussed in related work [12], this approach can produce gradients that are near-0 in some regions and near-infinite in others. This is especially true with rigid contact dynamics, longer trajectories, and simulations with a small timestep $\Delta t$.
@@
Approximately minimizing graph distance is the essence of the violation-implicit loss, and has been used in geometry learning given a known object trajectory [11]. The analogous procedure in our setting is to bring the near-discontinuous dynamics and physical constraints into the loss function as penalties, and the inner optimization happens over the contact impulses $\lambda$.
@@
Without dynamics as a hard constraint, the entire trajectory $x_{t}$ can be treated as decision variables, reducing the instability of differentiating through a large $T$ at the cost of introducing non-physical local minima in the optimization. In practice, we mitigate the latter using multiple trajectory initializations for the gradient descent. Following [11], our physical loss terms $g_{\theta}$ include the following.
@@
While [11] was agnostic to elasticity given the known trajectory, since we are estimating an unknown trajectory, we found it useful to assume that all of our collisions would be nearly inelastic. With the above loss terms, our inner optimization loop in $\lambda$ was a quadratic program (QP) with a second-order cone constraint that could be solved quickly in Cvxpylayers [35]. Note that, by the envelope theorem, explicitly differentiating through the $\min_{\lambda}$ is unnecessary, further speeding up the backwards pass.
@@ V Maximizing Expected Information Gain (EIG)
In Sec. V-A, we first review EIG as a popular metric for general active exploration [36, 14, 15]. In Sec. V-B, we detail the specific formulation for our setting: active tactile exploration with a dynamic unknown object and a violation-implicit loss (Sec. IV).
@@ V-A Preliminaries
For a Gaussian, the entropy can be expressed in terms of the log-determinant of the covariance: $H[\mathcal{N}(\mu,\Sigma)]=1/2\log\det\Sigma+const$. This is why it is common to use the (co)variance of the belief as a heuristic for high-information regions. But for dynamic objects, propagating a Gaussian prior through time to get that covariance is difficult. The result is fundamentally multi-modal, and just fitting a Gaussian can be completely non-physical. However, it is possible to fit a Gaussian to a second-order approximation of an arbitrary distribution about any modal peak. In that case, the effective covariance is equal to the Hessian of the log of the distribution about that peak ([36], Prop. 3.5).
@@
For a full derivation we direct the reader to [36], Sec. 3-5.
@@ V-B Computing EIG
While the gradients can be evaluated at the MLE trajectory $\tilde{x}_{0\cdots T}$, they must be taken with respect to the parameters we actually care about: $\tilde{\theta},\tilde{x}_{T}$. Frustratingly, this necessitates quantifying the sensitivity of past measurements to the current state. Even with a differentiable simulation, contact dynamics cannot in general be simulated backwards (multiple starting states can lead to the same end state). One could consider taking the dynamics Jacobian $\frac{dx_{t}}{dx_{t-1}}$ and directly inverting it to try and approximate this sensitivity, but in our experience this leads to severe numerical instability. For this work, we found success computing the sensitivity at each time step with respect to only that time step’s state and summing the results, e.g., $\sum_{t}\nabla_{x_{t}}\phi_{t}\nabla_{x_{t}}\phi_{t}^{T}$. This is equivalent to assuming $\frac{dx_{t}}{dx_{t-1}}=\mathbf{I}$. While sufficient for the sliding object trajectories that emerged in our experiments, we do not believe this approach is likely to work for more chaotic disturbances, which is an avenue for future work.
@@
The computation is identical to Eq. 14, with the approximation becoming equality as Fisher information is taken in expectation. The same note from the previous paragraph applies for gradient computation with respect to $\tilde{x}_{T}$. Differentiable calculation of $\phi$ and $\hat{n}$ (itself a QP) were implemented with Jaxopt [37]. With both Fisher and observed information, we can directly compute EIG with Eq. 11.
@@ VI-A Action Space
For the computation of bCH for evaluation only, FoundationPose [17] visually computes the ground-truth object pose at 30Hz. As we do not have any size prior for the object, we ignore the initial uniform sweep of the workspace and only begin counting actions at the moment of first contact. In practice, this means randomly selecting an action for one finger centered on the ground truth position.
Figure 3: Hardware setup: a Trifinger [38] modified to have 2 fingers directly opposing each other, each instrumented with a Densetact [3] sensor. Densetact images (green) can be used to locate the center-of-pressure (yellow), which in turn can be projected onto the nominal model to compute the measured normal vector $\hat{n}_{m,t}$ (pink). A single RealSense running FoundationPose (red) is used for evaluation only to calculate bCH.
@@ VI-B Hardware Setup
Our hardware setup is summarized in Fig. 3. We utilize a Trifinger robot [38] modified into a dual-finger configuration 180-degrees apart. Each finger is instrumented with a Densetact [3], which streams sensor images at 30Hz. Tactile data acquisition leveraged the correlation between observed fingertip motion primitives from each frame and applied forces (normal, shear, and torsional). Contact motion was computed using the optical flow algorithm [39] between the current frame and a reference frame with no contact. We perform a Helmholtz-Hodge decomposition [40] of the resulting vector field to separate out the irrotational component. Empirically, we find the surface deformation to be correlated with the scalar potential $\psi(u,v)$ of this field. The center-of-pressure (CoP) can be measured by taking a weighted average $\sum_{u,v}\psi(u,v)\cdot(u,v)$, and the contact normal can be imputed from a model of the sensor’s surface. For the contact boolean, we place a threshold on $\psi$ near the CoP.
@@ VII-A Simulation Results
We first evaluated our method in a Drake [41] simulation running at 1kHz (with data still reported at 30Hz). For the cuboid parameterization, our ground truth was a box with a randomized shape between 3 and 6cm, with our initial guess randomly sized between 1 and 3cm (in practice, a smaller initial guess lead to more stable learning). For the polytope parameterization, we used Platonic solids of fixed size (2cm bounding radius) as the ground-truth, and our initial guess was always 20 random vertices sampled from the 1cm sphere. Each trial, we initialized both the ground truth and our initial guess to a random pose within 2cm of the center of the workspace. After the initial random action that made contact with the ground truth, we took 5 (for cuboid) or 8 (for polytope) subsequent actions, recording bCH after each one. The results are summarized in Fig. 4.
@@
For the polytope parameterization, Fig. 6 (middle) shows 5 actions (excluding actions 3, 6, and 7 for brevity) of an example 8-action learning trajectory. Both EIG and Random exploration reach a bCH of 5mm consistently within 8 actions (8s of data). While not rising to the same statistical significance, a clear difference can still be seen between EIG and Random, with the latter requiring 3 extra actions for the 95% confidence interval to drop below 1cm. Note that this interval for Random already omits a catastrophic outlier, where the centroid of the guess was completely outside of the ground truth object, leading to 2 actions without contact before they re-aligned. Combined with an observable reduction in the variance of bCH across all actions, we can confidently say that the data supports the claim that EIG-maximizing exploration provides a more consistent learning trajectory. At the same time, even Random exploration still achieved consistent results with 5-8s of tactile data.

### v2

@@ I Introduction
Robot manipulators in the wild will inevitably come across previously unseen objects and environments. They can benefit from building object models from the limited data available at test time. Such models can be reused across different tasks and systems, and they can make learned manipulation policies more interpretable. In the building of these models, tactile sensing can compliment visual modalities like RGB and depth. Tactile data is robust to darkness, reflections, and occlusions from clutter, the environment, and the object itself. With the advent of inexpensive sensors [lambeta_digit_2020, johnson_microgeometry_2011, do_densetact_2023], touch is becoming a more common component of robot manipulation systems [weinberg_survey_2024].
Two challenges inherent to tactile sensing are sparsity and disturbance. By definition, data is only collected about the part of the object in direct physical contact with the sensor, necessitating multiple movements to get a complete picture. Sparsity can be addressed by active exploration, where actions are taken to efficiently maximize some information metric, often captured by maintaining a belief distribution over possible object parameterizations [hejna_robot_2025, schneider_active_2022, yi_active_2016, driess_active_2017, zhang_active_2017, shahidzadeh_actexplore_2024]. However, reasoning about information gain is complicated by disturbance: every contact has the potential to move the object unless it is heavy or bolted down. Propagating belief through motion is tricky, informing our first key insight: expected information gain (EIG), by foregoing an explicit belief distribution, is a particularly useful metric for active tactile exploration.
Disturbance affects even the model-building itself, as it necessitates learning the object’s pose trajectory in addition to its geometry. That trajectory needs to adhere to the physical constraints of rigid-body motion and contact, which can lead to difficult loss landscapes with both nearly flat and nearly discontinuous regions. Related work in geometry-learning with a known trajectory [pfrommer_contactnets_2020, bianchini_generalization_2022] has shown how a violation-implicit loss leads to a landscape more amenable to gradient-based optimization. Our second key insight is that this loss works just as well with the trajectory itself as a decision variable.
@@ II-A Model Building and System Identification
Visual modeling of both environment and object geometry is a mature field with the well-studied paradigms of simultaneous localization and mapping (SLAM) and structure from motion (SfM)[ozyesil_survey_2017]. Continuous object representations include radiance fields [jiang_fisherrf_2024] and Gaussian splatting [jiang_ag-slam_2024]. Recent work learns the geometry [wen_bundlesdf_2023] and pose [wen_foundationpose_2024] of dynamic objects in video, even in contact-rich environments [bianchini_vysics_2025]. For assessing the accuracy of our tactile-only system, we compare against ground truth estimates from FoundationPose [wen_foundationpose_2024].
Even without active exploration, tactile sensing has been used in recent work to refine models of both the robot and its environment. Non-optical contact sensors can enable more accurate proprioception in the presence of a known environment [koval_pose_2015]. Optical tactile sensors create estimates of local surface geometry that can be used to localize the robot with respect to a known object [suresh_midastouch_2023] or vice versa [sodhi_patchgraph_2022].
@@ II-B Active Exploration in Robotics
Active exploration has been studied in robotics for decades [bajcsy_active_1988] and involves either the selection of data-gathering actions to maximize information or the curation of data based on information potential. Generally, it includes: (1) some current state of knowledge, (2) an estimate of the information collected by a given action, and (3) an optimization metric quantifying how useful the latter is given the former.
In Bayesian state estimation, the current state can be represented by a belief distribution. Mutual information (MI) [hejna_robot_2025] as a metric captures the information gained about this belief distribution by observing some other (e.g. measurement likelihood) distribution, but in dynamic environments this can be intractable to compute exactly [schneider_active_2022]. Explicit MI calculation can be avoided by using uncertainty instead, e.g., probing areas where the belief distribution has a high variance [yi_active_2016, driess_active_2017]. One could also use a heuristic to avoid actions similar to those previously taken [shahidzadeh_actexplore_2024] or a policy trained to take informative actions [zhang_active_2017]. For dynamic objects, propagating belief though a dynamics model that is itself uncertain is tricky: the result is inherently multimodal, and the number of required parameters scales exponentially in the number of moments. Multimodal uncertainty can be approximated by simulating an ensemble of possible models [pathak_self-supervised_2019].
Another approach is to drop the belief distribution and maximize Fisher Information [jiang_fisherrf_2024], which only requires a point maximum-likelihood estimate (MLE) and is used extensively in optimal experimental design. Recent work uses this approach to refine a simulator from real world data [memmel_asid_2024] and plan informative trajectories in contact-rich environments [sathyanarayan_behavior_2025]. However, maximizing Fisher information alone does not take into account previously observed information over multiple actions. Ergodic exploration [abraham_data-driven_2018, abraham_ergodic_2017] can be used to plan a sequence of actions such that, while more time is spent in areas likely to have high information, the entire state space is eventually explored. EIG (see Sec. V), derived from Fisher information, explicitly quantifies how much information an action is likely to obtain relative to previously observed information. It has been used in visual model building for active view selection of static objects [jiang_fisherrf_2024, jiang_ag-slam_2024]. This work takes this procedure and implements it for tactile sensing of dynamic objects.
@@ II-C Tactile Active Exploration
Multiple works have tackled the problem of tactile active exploration. The most common application is the modeling of an unknown surface using poking or sliding contact from an optical sensor [hu_active_2024, ketchum_active_2024, yi_active_2016], as thoroughly collated by a recent survey [schneider_tactile_2025]. It is also common to estimate the geometry of 3d static objects [shahidzadeh_actexplore_2024, driess_active_2017]. When learning both pose and geometry, the latter tends to be treated as a classification problem by maintaining a discrete set of possible object geometries [kim_tactile_2024, xu_tandem3d_2023, zhang_active_2017]. One work simultaneously estimates the pose and geometry of specifically planar objects [suresh_tactile_2021]. To our knowledge, ours is the first work that learns the pose and geometry (from a continuous class of convex shapes) of a 3D rigid body that is free to move during deployment.
@@ IV Learning with a Violation-Implicit Loss
Where $f_{\theta}$ are the discrete-time dynamics as a function of contact impulses $\lambda_{t}$. These can themselves be written as the minimizer of a function $g_{\theta}$ [anitescu_optimization-based_2006] representing other physical constraints (e.g. non-penetration) discussed in more detail below. Under the assumption of Coulomb friction, impulses are constrained to the friction cone $\mathcal{FC}=\{\|\lambda_{t}^{f}\|_{2}\leq\mu\lambda_{t}^{n}\}$, where the total frictional impulse $\|\lambda_{t}^{f}\|_{2}$ is bounded by a constant multiple of the normal impulse $\lambda_{t}^{n}$, sliding in equality and sticking otherwise.
One could optimize this loss function through shooting and differentiable simulation. With parameter $x_{0}$, compute $x_{t}=(f_{\theta})^{t}(x_{0})$ and perform gradient descent through $(f_{\theta})^{t}$ to find $\tilde{x}_{0}$. Then $\tilde{x}_{T}=(f_{\theta})^{T}(\tilde{x}_{0})$. As discussed in related work [bianchini_generalization_2022], this approach can produce gradients that are near-0 in some regions and near-infinite in others. This is especially true with rigid contact dynamics, longer trajectories, and simulations with a small timestep $\Delta t$.
@@
Approximately minimizing graph distance is the essence of the violation-implicit loss, and has been used in geometry learning given a known object trajectory [pfrommer_contactnets_2020]. The analogous procedure in our setting is to bring the near-discontinuous dynamics and physical constraints into the loss function as penalties, and the inner optimization happens over the contact impulses $\lambda$.
@@
Without dynamics as a hard constraint, the entire trajectory $x_{t}$ can be treated as decision variables, reducing the instability of differentiating through a large $T$ at the cost of introducing non-physical local minima in the optimization. In practice, we mitigate the latter using multiple trajectory initializations for the gradient descent. Following [pfrommer_contactnets_2020], our physical loss terms $g_{\theta}$ include the following.
@@
While [pfrommer_contactnets_2020] was agnostic to elasticity given the known trajectory, since we are estimating an unknown trajectory, we found it useful to assume that all of our collisions would be nearly inelastic. Without this term, the object could spontaneously gain energy from the ground contact force, bouncing non-physically. With the above loss terms, our inner optimization loop in $\lambda$ was a quadratic program (QP) with a second-order cone constraint that could be solved quickly in Cvxpylayers [cvxpylayers2019]. Note that, by the envelope theorem, explicitly differentiating through the $\min_{\lambda}$ is unnecessary, further speeding up the backwards pass.
@@ V Maximizing Expected Information Gain (EIG)
In Sec. V-A, we first review EIG as a popular metric for general active exploration [kirsch_unifying_2022, jiang_fisherrf_2024, jiang_ag-slam_2024]. In Sec. V-B, we detail the specific formulation for our setting: active tactile exploration with a dynamic unknown object and a violation-implicit loss (Sec. IV).
@@ V-A Preliminaries
For a Gaussian, the entropy can be expressed in terms of the log-determinant of the covariance: $H[\mathcal{N}(\mu,\Sigma)]=1/2\log\det\Sigma+const$. This is why it is common to use the (co)variance of the belief as a heuristic for high-information regions. But for dynamic objects, propagating a Gaussian prior through time to get that covariance is difficult. The result is fundamentally multi-modal, and just fitting a Gaussian can be completely non-physical. However, it is possible to fit a Gaussian to a second-order approximation of an arbitrary distribution about any modal peak. In that case, the effective covariance is equal to the Hessian of the log of the distribution about that peak ([kirsch_unifying_2022], Prop. 3.5).
@@
For a full derivation we direct the reader to [kirsch_unifying_2022], Sec. 3-5.
@@ V-B Computing EIG
While the gradients $\nabla_{\omega}$ can be evaluated at the MLE trajectory $\tilde{x}_{0\cdots T}$, they must be taken with respect to the parameters we actually care about: $\omega=\{\tilde{\theta},\tilde{x}_{T}\}$. Frustratingly, this necessitates quantifying the sensitivity of past measurements to the current state, as if through a backward simulation $x_{T-1}=f^{-1}(x_{T})$. Even with a differentiable simulation, contact dynamics cannot in general be simulated backwards (multiple starting states can lead to the same end state). One could consider taking the dynamics Jacobian $\frac{dx_{t}}{dx_{t-1}}$ and directly inverting it to try and approximate this sensitivity, but in our experience this leads to severe numerical instability. For this work, we found success computing the sensitivity at each time step with respect to only that time step’s state and summing the results, e.g., $\sum_{t}\nabla_{x_{t}}\phi_{t}\nabla_{x_{t}}\phi_{t}^{T}$. This is equivalent to assuming $\frac{dx_{t}}{dx_{t-1}}=\mathbf{I}$. While sufficient for the sliding object trajectories that emerged in our experiments, we do not believe this approach is likely to work for more chaotic disturbances, such as if the object tumbles over. This is an active avenue for future work.
@@
The computation is identical to Eq. 14, with the approximation becoming equality as Fisher information is taken in expectation. The same note from the previous paragraph applies for gradient computation with respect to $\tilde{x}_{T}$. Differentiable calculation of $\phi$ and $\hat{n}$ (itself a QP) were implemented with Jaxopt [jaxopt_implicit_diff]. With both Fisher and observed information, we can directly compute EIG with Eq. 11.
@@ VI-A Action Space
For the computation of bCH for evaluation only, FoundationPose [wen_foundationpose_2024] visually computes the ground-truth object pose at 30Hz. As we do not have any size prior for the object, we ignore the initial uniform sweep of the workspace and only begin counting actions at the moment of first contact. In practice, this means randomly selecting an action for one finger centered on the ground truth position.
Figure 3: Hardware setup: a Trifinger [wuthrich_trifinger_2021] modified to have 2 fingers directly opposing each other, each instrumented with a Densetact [do_densetact_2023] sensor. Densetact images (green) can be used to locate the center-of-pressure (yellow), which in turn can be projected onto the nominal model to compute the measured normal vector $\hat{n}_{m,t}$ (pink). A single RealSense running FoundationPose (red) is used for evaluation only to calculate bCH.
@@ VI-B Hardware Setup
Our hardware setup is summarized in Fig. 3. We utilize a Trifinger robot [wuthrich_trifinger_2021] modified into a dual-finger configuration 180-degrees apart. Each finger is instrumented with a Densetact [do_densetact_2023], which streams sensor images at 30Hz. Tactile data acquisition leveraged the correlation between observed fingertip motion primitives from each frame and applied forces (normal, shear, and torsional). Contact motion was computed using the optical flow algorithm [teed_raft_2020] between the current frame and a reference frame with no contact. We perform a Helmholtz-Hodge decomposition [tactile_hhd] of the resulting vector field to separate out the irrotational component. Empirically, we find the surface deformation to be correlated with the scalar potential $\psi(u,v)$ of this field. The center-of-pressure (CoP) can be measured by taking a weighted average $\sum_{u,v}\psi(u,v)\cdot(u,v)$, and the contact normal can be imputed from a model of the sensor’s surface. For the contact boolean, we place a threshold on $\psi$ near the CoP.
@@ VII-A Simulation Results
We first evaluated our method in a Drake [tedrake_manipulation] simulation running at 1kHz (with data still reported at 30Hz). For the cuboid parameterization, our ground truth was a box with a randomized shape between 3 and 6cm, with our initial guess randomly sized between 1 and 3cm (in practice, a smaller initial guess lead to more stable learning). For the polytope parameterization, we used Platonic solids of fixed size (2cm bounding radius) as the ground-truth, and our initial guess was always 20 random vertices sampled from the 1cm sphere. Each trial, we initialized both the ground truth and our initial guess to a random pose within 2cm of the center of the workspace. After the initial random action that made contact with the ground truth, we took 5 (for cuboid) or 8 (for polytope) subsequent actions, recording bCH after each one. The results are summarized in Fig. 4.
@@
For the polytope parameterization, Fig. 6 (middle) shows 5 actions (excluding actions 3, 6, and 7 for brevity) of an example 8-action learning trajectory. Both EIG and Random exploration reach a bCH of 5mm consistently within 8 actions (8s of data). While not rising to the same statistical significance, a clear difference can still be seen between EIG and Random, with the latter requiring 3 extra actions for the 95% confidence interval to drop below 1cm. Note that this interval for Random already omits a catastrophic outlier, where the centroid of the guess was completely outside of the ground truth object, leading to 2 actions without contact before they re-aligned. This is an example of a suboptimal local minimum arising from the learning scheme. Combined with an observable reduction in the variance of bCH across all actions, we can confidently say that the data supports the claim that EIG-maximizing exploration provides a more consistent learning trajectory. At the same time, even Random exploration still achieved consistent results with 5-8s of tactile data.